OpenStudy (anonymous):

Hi I got a problem 'bout Fourier series, my teacher said to find Ao = 1/L int f(x) from -L to L but in many literatures I found to Find Ao = 1/2L int f(x) from -L to L... which one is right? thanks alot

OpenStudy (anonymous):

Some texts will give you the Fourier series as \(\dfrac{1}{2}a_0+\sum\cdots\), where \(a_0=\dfrac{1}{L}\int_{-L}^Lf(x)~dx\). Others will incorporate the 1/2 into the integral formula for \(a_0\). Same difference; just make sure you remember, if you use one of the formulas, to not use the other at the same time.

OpenStudy (anonymous):

oh, okay Thanks a lot.. :)